OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!!!!!! The area of a rectangular piece of land is 260 square meters. If the length of the land was 5 meters less and the width was 2 meters more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. Part B: What is the width of the land in meters? Show the steps of your work.

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@iambatman @nincompoop

OpenStudy (agent0smith):

Call the length L and the width W. Area is 260: L*W = 260 If it was a square, the length and width would be equal... If the length of the land was 5 meters less and the width was 2 meters more,: L-5 = W+2 You now have two equations and two unknowns: L*W = 260 L-5 = W+2 Solve, using substitution.

OpenStudy (anonymous):

can you show me how to do that please?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

@agent0smith ?